OpenStudy (anonymous):

Which of the following is a solution to the system of linear equations below? 2x=-3y-11 2y=6x-22

OpenStudy (anonymous):

A. (2, 5) B. (–2, 5) C. (2, –5) D. (–2, –5)

OpenStudy (dangerousjesse):

Solve the following system: 2 x = -11-3 y 2 y = 6 x-22 Express the system in standard form: 2 x+3 y = -11 -(6 x)+2 y = -22 Swap equation 1 with equation 2: -(6 x)+2 y = -22 2 x+3 y = -11 Add \(\large \frac{1}{3} ×\) (equation 1) to equation 2: \[\large -(6 x)+2 y = -22 \]\[\large 0 x+\frac{11 y}{3} = -\frac{55}{3}\]Divide equation 1 by 2: -(3 x)+y = -11 0 x+(11 y)/3 = -55/3 Multiply equation 2 by \(\large \frac{3}{11}\): \[\large -(3 x)+y = -11 \]\[\large 0 x+y = -5 \]Subtract equation 2 from equation 1: \[\large -(3 x)+0 y = -6 \]\[\large 0 x+y = -5 \]Divide equation 1 by -3: \[\large x+0 y = 2 \]\[\large 0 x+y = -5 \]Collect results:

OpenStudy (midhun.madhu1987):

 now find x by substituting the value of y in any equation...